OpenStudy (anonymous):

Find an equation in standard form for the ellipse with the vertical major axis of length 10, and minor axis of length 8.

OpenStudy (zale101):

Standard form of ellipse=\[y^2/a^2 +x^2/b^2=1\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Lol, this is a good questions. The equation of a circle is x^2 + y^2 = z^2 so if memory serves me, an ellipse is just a circle with different factors a and b dividing the variables.

OpenStudy (anonymous):

wait soooo?

OpenStudy (anonymous):

ahh wait I think I get it

OpenStudy (zale101):

the equation i put above was supposed to be the standard for of an vertical ellipse. I think that's what your question meant. major axis=2b so b=5 minor axis =2a so a=4

OpenStudy (anonymous):

ANSWER: